A ship travels 18.5 km on a course of 31.2 degrees south of east, then 12.7 km due south, then 21.5 km on the course of 61.3 degrees west of south, where it lands. Find the displacement and the angle traveled from the starting point to the landing point.
1
Expert's answer
2011-06-22T10:15:34-0400
Let’s find projections of the vector of displacement.
Vx = 18.5*cos31.2 - 21.5*sin61.3 ≈ -3.0344
Vy = 18.5*sin31.2 + 12.7 + 21.5*cos61.3 ≈ 32.6083

Displacement = √(Vx2+Vy2 ) = 32.7492
Angle = 90 - arctan((|Vy|)/(|Vx|)) ≈ 5.333 degrees to the west of south.